OpenStudy (brro2826):

Are these equations similar or are they different? x(900-x/2) and -1/2 x^2 + 450x I think they are similar but do not understand where the -1/2 would come from.

OpenStudy (mathmale):

Multiply out the first expression (it's not an equation, by the way). Then compare your result to the 2nd expression.

OpenStudy (sshayer):

\[x \left( \frac{ 900-x }{ 2 } \right)=\frac{ 900x-x^2 }{ 2 }=\frac{ 900x }{ 2 }-\frac{ 1 }{ 2 }x^2\]

OpenStudy (mathmale):

You could simplify this by multiplying ALL terms by 2. This eliminates fractions. Do that. Are the left and right hand expressions equal, or not?